# Drawing Pandas: A Benchmark for LLMs in Generating Plotting Code

### v1

@@ Abstract
This paper introduces the human-curated PandasPlotBench dataset, designed to evaluate language models’ effectiveness as assistants in visual data exploration. Our benchmark focuses on generating code for visualizing tabular data—such as a Pandas DataFrame—based on natural language instructions, complementing current evaluation tools and expanding their scope. The dataset includes 175 unique tasks. Our experiments assess several leading Large Language Models (LLMs) across three visualization libraries: Matplotlib, Seaborn, and Plotly. We show that the shortening of tasks has a minimal effect on plotting capabilities, allowing for the user interface that accommodates concise user input without sacrificing functionality or accuracy. Another of our findings reveals that while LLMs perform well with popular libraries like Matplotlib and Seaborn, challenges persist with Plotly, highlighting areas for improvement. We hope that the modular design of our benchmark will broaden the current studies on generating visualizations. Our benchmark is available online: https://huggingface.co/datasets/JetBrains-Research/plot_bench. The code for running the benchmark is also available: https://github.com/JetBrains-Research/PandasPlotBench.
@@ I Introduction
The domain of data visualization benchmarks for LLMs encompasses various tasks, including generating concise Matplotlib code snippets [6], visually guided code generation [4, 7, 8], as well as data analysis tasks leveraging SQL queries and visualization languages [9]. However, only one work, MatPlotBench [4], addresses the task of visualizing data from a given DataFrame or a CSV file, although it comprises only a limited number of data points. Moreover, current benchmarks fall short in providing an in-depth examination of how rephrasing instructions and utilizing different plotting libraries can impact performance, as well as benchmarking across a wide range of models.
@@
These varied results represent an important first step in a detailed study of modern LLMs in generating plotting code. While they are important in themselves, we highlight them to showcase the deep contents of PandasPlotBench, pointing the way to how it can be used to extend the field in the future.
@@ II-C Evaluation Metrics
Using the provided model, our benchmark generates the plotting code according to the given task. This code is then executed within a Jupyter Notebook environment, which is subsequently parsed to obtain the plot image. Each plotting code is executed in an individual notebook cell. We track the ratio of incorrect code, defined as cells that did not result in a plot. When the notebook cells return both plots and errors, we ignore the errors. The scoring is performed by the multimodal GPT-4o Judge model in two ways:
Visual scoring. The Judge model is asked to compare the generated and the ground truth plots on 0–100 scale, focusing on the main idea of the plots while ignoring styling details.
@@
It is important to note that the generated images do not have to match the ground-truth plots exactly, as the tasks were provided as text and are not extremely detailed. Additionally, the dataset tasks were generated by the GPT-4V model based on the given code and plot image. Therefore, some information may have been lost during this two-step process. In our opinion, the generated plots were often better than the ground-truth ones, leading us to include task-based scoring, which generally yields higher scores on average. Moreover, the Pearson correlation between task-based scores and visual scores for the best-performing model is 0.58, highlighting the importance of using both approaches. Judging prompts are provided in the supplementary materials [20] and are based on the corresponding prompts from the work of Yang et al. [4].
@@ III-B Benchmarking Plotting Libraries
To evaluate the model’s ability to generate code across different plotting libraries, we requested GPT-4o to produce code for Matplotlib, Seaborn, and Plotly (see Table II). We explicitly instructed the model to use the target library and not others. For the generated code, we checked whether the target library was imported. If the target library was not imported, the score was set to 0, and such errors were aggregated in the "Wrong library" column. Both Matplotlib and Seaborn libraries demonstrated high scores. Since the ground truth for visual scoring was always the Matplotlib score, the visual-guided scoring was predictably lower for other libraries. Still, Plotly was much less familiar to GPT-4o, resulting in 38 instances (22%) of non-functional code. We manually reviewed these cases and found that most errors were due to the incorrect usage of the Plotly library API by the language model.
@@ V Limitations and Future Work
The benchmark is based on open-source and synthetically generated data. Although the synthetic tasks and data files have never been seen by any LLMs, models may still exhibit bias towards publicly available ground-truth Matplotlib code. However, this code has been modified, and solutions using other libraries remain unseen. Another potential issue could be a bias towards OpenAI models, which were used for synthetic task generation, but we believe this bias to be negligible. Therefore, we argue that our tasks are less prone to models memorizing training data, as there is no direct link between the answers to benchmark tasks and the raw repository data that modern models use for training.
Our dataset is currently limited to 175 data points originating from the Matplotlib gallery. Future work should certainly include expansion to other sources. Because of this limitation, we used the Matplotlib ground truth plots for evaluating other libraries too, relying more on the task-based scoring, but it is important to collect ground truth data from other frameworks. Also, the DataFrames in the dataset are mostly concise, containing only the data required for plotting. Augmenting and modifying these files with random columns and additional data could make the generation task more representative of real-world scenarios and more complex.

### v2

@@ Abstract
This paper introduces the human-curated PandasPlotBench dataset, designed to evaluate language models’ effectiveness as assistants in visual data exploration. Our benchmark focuses on generating code for visualizing tabular data—such as a Pandas DataFrame—based on natural language instructions, complementing current evaluation tools and expanding their scope. The dataset includes 175 unique tasks. Our experiments assess several leading Large Language Models (LLMs) across three visualization libraries: Matplotlib, Seaborn, and Plotly. We show that the shortening of tasks has a minimal effect on plotting capabilities, allowing for the user interface that accommodates concise user input without sacrificing functionality or accuracy. Another of our findings reveals that while LLMs perform well with popular libraries like Matplotlib and Seaborn, challenges persist with Plotly, highlighting areas for improvement. We hope that the modular design of our benchmark will broaden the current studies on generating visualizations. Our dataset and benchmark code are available online: https://huggingface.co/datasets/JetBrains-Research/PandasPlotBench; https://github.com/JetBrains-Research/PandasPlotBench.
@@ I Introduction
The domain of data visualization benchmarks for LLMs encompasses various tasks, including generating concise Matplotlib code snippets [6], visually guided code generation [4, 7, 8], as well as data analysis tasks leveraging SQL queries and visualization languages [9]. However, only one work, MatPlotBench [4], addresses the task of visualizing data from a given DataFrame or a CSV file, although it comprises only a limited number of data points (25 datapoints using CSV data files). Moreover, current benchmarks fall short in providing an in-depth examination of how rephrasing instructions and utilizing different plotting libraries can impact performance, as well as benchmarking across a wide range of models.
@@ II-C Evaluation Metrics
Using the provided model, our benchmark generates the plotting code according to the given task. This code is then executed within a Jupyter Notebook environment, which is subsequently parsed to obtain the plot image. We chose Jupyter as a universal environment that enables seamless switching between plotting libraries and easy parsing of plot images. Each plotting code is executed in an individual notebook cell. We track the ratio of incorrect code, defined as cells that did not result in a plot. When the notebook cells return both plots and errors, we ignore the errors. The scoring is performed by the multimodal GPT-4o Judge model in two ways:
Visual scoring. The Judge model is asked to compare the generated and the ground truth plots on 0–100 scale, focusing on the main idea of the plots.
@@
It is important to note that the generated images do not have to match the ground-truth plots exactly, as the tasks were provided as text and are not extremely detailed. Additionally, the dataset tasks were generated by the GPT-4V model based on the given code and plot image. Therefore, some information may have been lost during this two-step process. In our opinion, the generated plots were often better than the ground-truth ones, leading us to include and prefer task-based scoring, which generally yields higher scores on average. Moreover, the Pearson correlation between task-based scores and visual scores for the best-performing model is 0.58, highlighting the importance of using both approaches. Judging prompts are provided in the supplementary materials [20] and are based on the corresponding prompts from the work of Yang et al. [4].
@@ III-B Benchmarking Plotting Libraries
To evaluate the model’s ability to generate code across different plotting libraries, we requested GPT-4o to produce code for Matplotlib, Seaborn, and Plotly (see Table II). We explicitly instructed the model to use the target library and not others. For the generated code, we checked whether the target library was imported. If the target library was not imported, the score was set to 0, and such errors were aggregated in the "Wrong library" column. Both Matplotlib and Seaborn libraries demonstrated high scores. Since the ground truth for visual scoring was always the Matplotlib score, the visual-guided scoring was predictably lower for other libraries. Still, Plotly was much less familiar to GPT-4o, resulting in 38 instances (22%) of incorrect code. We manually reviewed these cases and found that most errors were due to the incorrect usage of the Plotly library API by the language model.
@@ V Limitations and Future Work
The benchmark is based on open-source and synthetically generated plotting data. Although the synthetic tasks and data files have never been seen by any LLMs, models may still exhibit bias towards publicly available ground-truth Matplotlib code. However, this code has been modified, and solutions using other libraries remain unseen. Another potential issue could be a bias towards OpenAI models, which were used for synthetic task generation, but we believe this bias to be negligible. Therefore, we argue that our tasks are less prone to models memorizing training data, as there is no direct link between the answers to benchmark tasks and the raw repository data that modern models use for training.
Our dataset is currently limited to 175 data points originating from the Matplotlib gallery. Future work should include expansion to other sources. Due to this limitation, we used the Matplotlib ground truth plots to evaluate other libraries, relying more on the task-based scoring. Also, the DataFrames in the dataset are mostly concise, containing only the data required for plotting, often far from real-world data. Enriching, augmenting, and modifying these files with random columns and additional data could make the generation task more representative of real-world scenarios and more complex.